OpenStudy (anonymous):

Solve 6/x = 3/8 for x.

OpenStudy (karatechopper):

9/4=x

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

when two ratios are equal, then the cross products of the ratios are equal. that is, for the proportion, a:b = c:d , a x d = b x c

OpenStudy (anonymous):

2.25 is =x ?

OpenStudy (anonymous):

if 6/x=3/8, then 3x=48. so by dividing both sides 3, you get x=16.

OpenStudy (anonymous):

Cross-multiply